OpenStudy (anonymous):

Find all numbers c that satisfy the conclusion of Rolle's Theorem for the following function and interval. f(x)=8 x^3- 24 x^2 + 16 x + 8, [0,2]

OpenStudy (anonymous):

f'(x)=24*(x^2)-48*x+16 f"(x)=48*x-48 48*x-48=0 x=1 \[-\infty lex\]